# Photo-nuclear cross sections on Au

### v1

@@ Abstract
A method was developed for measuring photonuclear reactions concurrently at several discrete photon beam energies on a stack of different target materials via a single irradiation. Concentric ring targets of the materials (in order from front to back targets: Au, TiO, Zn, Os, and Au) were irradiated at the High Intensity Gamma-ray Source (HIS). As a proof of principle, we report the result of the cross section measurements from the front Au target. The excitation functions of the Au(,n)Au and Au(,3n)Au reactions were determined in the incident photon energy range of 13-31 MeV using quasi-monoenergetic photon beams provided at HIS. The cross sections of the combined ground state (2) and short-lived first isomeric state (m1, 5), and of the second isomeric state (m2, 12) in the Au production are obtained separately by subtracting the rays from the internal conversion of the second isomeric state. The excitation function of the second isomeric state via the photon-induced reaction Au(,n)Au was measured for the first time. By using the activation method rather than direct neutron counting, the exclusive cross sections for the (,n) and (,3n) reactions were determined. Comparing the yields from the front and back gold targets validates our ability to simulate the effect of photon scattering in the target stack and provides a method for assessing the systematic uncertainty of our technique.
@@ I Introduction
Radioisotopes play important roles in numerous fields ranging from medical treatments to national security and basic research. Some examples include investigations of structures and reactions involving atomic nuclei, Mossbauer spectroscopy, radio-thermoelectric generation and other nuclear batteries, nuclear device detection, and the mitigation of nuclear proliferation [1]. The reports from “Workshop on the Nation’s Needs for Isotopes: Present and Future”, “Isotopes for the Nation’s Future, A long range plan”, and “Meeting Isotope Needs and Capturing Opportunities for the Future” identify multiple isotopes in short supply [1, 2, 3]. Among these are high specific activity beta emitters such as Sc, Cu, As, and Re. These radioisotopes are of interest to the community as they have ideal nuclear properties for medical applications. Their availability is limited as they are difficult to produce in high specific activity using common production methods.
@@
Most radioisotopes are produced using either nuclear reactors or light-ion accelerators. High specific activity radioisotopes are very important in nearly all aspects of radiochemical work. However, most radionuclides produced in nuclear reactors have low specific activity due to the inherently limited production pathways, e.g., dominated by (n,) reactions. Specific activity refers to the radioactivity of the produced radionuclide relative to the total mass of the element of interest present in the sample. In contrast, accelerator-produced radionuclides are generally characterized by high specific activity, e.g., (p,xn) reactions are typically used because the produced isotope is a different element from the target material. Electron linear accelerators (LINACs) are unique sources of radioisotopes. Even though the basic technology has been around for decades, only recently have electron LINACs capable of producing photons with sufficient energy and flux for radioisotope production become available. The Nuclear Science Advisory Committee (NSAC) recently named this approach as one of the most compelling with the largest-impact opportunities for the production of high specific activity radioisotopes [1]. Electron LINACs activate materials via photonuclear reactions. The yields of these reactions depend on the production of high-energy photons generated by interactions of high-energy electrons with a high-Z material (i.e., a converter) to produce Bremsstrahlung radiation. Production yields are controlled by the electron beam energy, beam flux, target size, length of irradiation, and reaction cross section. A great deal of work has been done to create experimental cross-section databases for (,n) reactions (leading towards low specific activity radioisotopes) [4, 5, 6]. However, these data are incomplete. For example, the cross-section data for (,p) reactions (which lead to high specific activity radioisotopes) still require considerable work. There are many theoretical databases available which provide guidance on photo-production of radionuclides. Experimental data are needed to assess the accuracy of predictions of the leading reaction models used in nuclear-data libraries (TENDL, JENDL, ENDF, JEFF, CENDL, BROND, etc.) [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Examples of theoretically predicted cross sections that incorrectly estimate the production rate of radioisotopes are available in the literature [17, 18, 19, 20]. One specific example is the photonuclear production of Cu via the reaction Zn(,p)Cu where the maximum theoretical cross section in the TENDL database is 3.7 mb at 22 MeV, but was experimentally determined to be 11.76 mb as reported by the Experimental Nuclear Reaction Data (EXFOR) database. This result effectively implies that Cu can be produced more economically via the (,p) reaction than previously thought. In order to better assess the feasibility of producing Cu and other radioisotopes with photonuclear reactions, (,n) or (,p), excitation functions must be experimentally evaluated. Direct measurement of photonuclear cross sections using electron beams is not trivial as converter-produced Bremsstrahlung photons deliver a beam with a broad energy spectrum to the targets [21, 22]. Compton back-scattering is an elegant solution to this problem. Compared with a Bremsstrahlung beam, with a broad-band energy continuum, a Compton back-scattered -ray beam is narrowly peaked around a specific energy. Very precise measurement of cross sections can be obtained using activation techniques with target irradiation by these quasi-monochromatic gamma beams. Such quasi-monoenergetic photon beams are available at the Triangle Universities Nuclear Laboratory’s (TUNL) High-Intensity Gamma-Ray Source (HIS) at the Duke Free-Electron-Laser Laboratory (DFELL). Photons with selectable energy are produced using intracavity backscattering of free-electron laser (FEL) photons from electrons in the DFELL storage ring. Gamma-ray beams with photon energies of up to 100 MeV and total intensities of 10 - 10 /s can be produced [23]. High-energy photons are produced by colliding electron bunches that are synchronized so that laser photons from one bunch are reflected from a downstream mirror and then collide head-on with the second electron bunch in a magnetic-field-free scattering region. Photon beam energies and fluxes are calibrated with in-line detectors and activation of well-defined monitor foils, respectively. Additionally, the back-scattered gamma beam has a well-defined radial energy dependence. The central portion of the beam has the highest energy. The beam energy and flux density decrease radially outward from the center in a predictable manner. The beam flux also surrenders to this phenomenon [24]. Activation of concentric ring targets by an uncollimated beam will provide experimental activation data that may be used to calculate cross sections at the energy of the beam incident on each ring. This method was developed and used for the first time at HIS in the present energy-dependent cross-section measurements which provide activation data at 3 energies for each target irradiation.
@@
In this paper, we describe our newly developed experimental techniques for measuring photonuclear cross sections using the laser-Compton scattered beam at HIS, and report cross sections for photon-induced reactions on Au measured using these new techniques. The cross section of Au(,xn) reactions are important for beam flux measurements in cross section experiments. For example, gold monitor foils are routinely used to normalize Bremsstrahlung photon beams in cross-section measurements [25, 26]. The Au(,n)Au cross section is very well characterized in the energy range of 8-20 MeV, which is below the threshold for multiple neutron emission (see Table 1). However, there are discrepancies between existing experimental data [27, 28], and above 20 MeV gamma energy, there are very few experimental data and theoretical predictions by reaction models differ substantially. These photon energies are above the threshold for reactions leading to high specific activity radionuclides via the (,p) reaction, such as Ti(,p)Sc and Zn(,p)Cu at around 20 MeV and higher. A goal of the present work is to measure the exclusive cross-section excitation functions of the Au(,n) and (,3n) reactions with improved accuracy over existing data for use as a standard in determining -ray beam fluxes. Table 1 provides a summary of the world data for the Au(,xn) reactions by beam type. These data are compared to the results of our new measurements in Section V. The results of this work for the Ti, Zn, and Os targets measured concurrently with Au will be presented in a subsequent paper. In this paper, our multiple target multi-energy technique is described in section II, the HIS Laser Compton Scattering (LCS) beam is described in section III, the experiment is discussed in section IV, section V describes the cross-section analysis, and the results are presented and discussed in section VI.
@@
| (,n) | O. Itoh et al. [32] | 2011 | 8.08-13.14 | 30 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,n) | F. Kitatani et al. [33] | 2011 | 9.3-12.4 | 12 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,n) | F. Kitatani et al. [34] | 2010 | 10.0-14.6 | 8 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,n) | K.Y. Hara et al. [35] | 2007 | 8.23-12.3 | 24 | AIST, Tsukuba, Japan | LCS source | neutron counting |
@@ II Method
This section details the method applied in this work to determine the exclusive (,n) and (,3n) cross sections via the activation method. The excitation functions reported here cover the gamma-energy range from 13 to 31 MeV. The photonuclear cross sections over the range of the Giant Dipole Resonance (GDR) vary by orders of magnitude. The goal of this work is to demonstrate our activation techniques by applying them in measuring the excitation functions of Au(,xn) reactions with a resolution of 3-10 at each energy. At HIS, the beam energy spread depends on the beam collimation. In a typical setup with an energy spread (FWHM) of 5, the flux transmitted through the collimator is about 7.5 of the total flux of the Compton-scattered laser beam. Our method of using 3 radial target segments enables 3 gamma energies of the excitation function to be measured concurrently. The radial energy and flux distribution of the HIS beam was verified experimentally by blocking portions of the beam with tungsten plugs to mimic the concentric-rings target design. Simulations were performed and compared to the experimental results of these calibrations to determine the accuracy of the beam simulations.
@@ II.1 Target stack
In the activation method, the photon beam induces radioactivity via photon-induced reactions, such as (,n) and (,p), in targets. Reaction yields are determined by gamma counting the targets after activation. The target foils used in these experiments are made of separable concentric rings that exploit the energy spread of the HIS beam and enable the activation of multiple targets across multiple energies in a single beam exposure. A concept of a target stack is illustrated in Figure 1. The relatively low and well characterized attenuation of high-energy photons by the target materials provides the ability to irradiate multiple targets in a single irradiation. The total thickness of the stack of targets used in these experiments was 2.8-3.3 g/cm, which resulted in a 13 total attenuation of the photon beam at 14 MeV. The entire target may then be exposed to the beam and each individual concentric ring will interact with a narrow range of beam energies and a calculable fraction of the overall photon flux. The individual rings are analyzed with high-purity germanium detectors (HPGe) to determine activation yields. Thus, using this technique, the cross sections at 3 energies can be measured in a single irradiation. With the stack of 5 targets and 3 segments per target, yields for 15 target segments are obtained for each 12-hour run.
@@ II.2 Activation
The reaction rate (R) is given by,
where is the cross section, is the number of atoms in the target, and I is the beam flux of incident particles.
An example of a radioisotope production curve is shown in Figure 2. The isotope is being produced faster than it can decay during the time from 0 to when the target is being irradiated.
II.2.1 Au(,n)Au
The Au nucleus has three different long-lived states; e.g. the ground, and isomeric states denoted by g (6.17 d), (8.1 s) and (9.6 h). The two isomeric states decay 100 to the ground state. The number of nuclei produced during the irradiation time can be calculated using the formulae below,
where the R’s and ’s are the reaction rates and decay constants, respectively. is the number of nuclei in the th state, here numbers 0,1 and 2 represent ground, isomeric and isomeric state, respectively. The solutions for the number of nuclei in the and isomeric states are obtained from the following equations.
The number of nuclei can be calculated from integrating equation 3,
@@ II.2.1 Au(,n)Au
The total number of events with measurement time() will be,
Using these results and the relation between the number of events and the cross section in equation 1, the cross section for producing Au in the 2 isomer state is,
with,
II.2.2
@@
The photon beams are produced by Compton scattering of FEL photons from relativistic electrons circulating in a storage ring at HIS. In the case of relativistic electrons ( 1) scattering at a small angle (1), the scattered photon energy after collision is given by,
where E is the scattered photon energy, and E and E are the energies of the incident photon and electron. The Lorentz factor of the electron is denoted as . The angle is expressed in terms of the radius of the target (r) and the distance from the collision point to the target position, r/L. The spatial distribution of the scattered photons can be obtained theoretically [24]. The theoretical predictions are verified using a computer code [24] developed by HIS’s accelerator group, which was benchmarked with other codes such as [36] and using experimental results at HIS [24]. As an example of Monte-Carlo simulation results based on the theoretical prediction, if we consider the -ray beam at HIS with E 660 MeV and 450 nm, the scattered -ray beam energies at the target location (L 53 m) at each of the three segments will be 18, 17 and 16 MeV, respectively. Typical percentile energy spreads (1) of the photon beam at the 1, 2 and 3 segments are 1.3-2.7, 2.2-4.3 and 3.2-5.7 of the central beam energy from 14 to 31 MeV, respectively. The spatial distributions and photon energies in each target segment from simulations are provided in Figure 3, respectively. The accuracy of the simulations of the beam energy profile in different rings concentric with the beam axis have been validated by measurements. These measurements were performed using rays produced under conditions identical to those used in the simulations, E 725 MeV and 350 nm. The energy spectrum of the beam was measured using a large cylindrical NaI detector (10” dia. 14” thick) positioned with its center on the -ray beam axis. This detector is referred to as “Molly”. Data were taken for four collimator arrangements: N1 = 0.0 to 6.35 mm; N2 = 0.0 to 15.9 mm; N3 = 6.35 to 15.9 mm; and N4 = 11.1 to 15.9 mm. The efficiency of the large NaI detector (Molly) is greater than 97 at the energies of these measurements. The N1 and N2 measurements were performed using standard lead collimators with openings of 0.5-inch and 1.25-inch diameter. The N3 and N4 measurements were carried out by inserting a 0.5-inch and a 0.875-inch diameter tungsten rod concentric with the 1.25-in diameter collimator, respectively. Both tungsten rods were 6 inches long.
The beam flux at each segment of each target is essential in determining the corresponding reaction cross sections. The fractional photon beam fluxes at each target segment were obtained from simulations. With these incident beam fluxes on the segments of the front target, the beam intensity at each target was calculated using the beam transport simulation performed by the PHITS Monte-Carlo code [37]. The geometry illustrated in Figure 4 was used for the simulations. The targets in the stack were Au, TiO, Zn, Os and Au. Successive targets were separated by 0.45 cm. The detailed information about the segments of the actual target stack can be found in Section IV. A total of 15 segments in the target stack (3 segments per target) were exposed to the photon beams at once. The beam flux was simulated for each irradiation (See Table 2). In this energy range, Compton scattering and pair production dominate the beam attenuation and scattering as it passes through the stack. The attenuation and scattering of the gamma beam is simulated in detail using the PHITS code with the overall attenuation predicted to be 10. The central beam energy, one-sigma radial energy spread, and beam flux on each segment of the stack are summarized in Table 2. The overall simulation is validated using the data from the gold target at the back of the stack when compared with the data from the gold target in the front.
| E | 14 MeV | 16 MeV | 18 MeV | 22 MeV | |||||||||
| seg | tgt. | flux () | flux () | flux () | flux () | ||||||||
| TiO | 13.956 | 0.182 | 23.080 | 15.944 | 0.226 | 23.430 | 17.906 | 0.243 | 23.840 | 21.825 | 0.497 | 23.536 | |
| TiO | 13.477 | 0.303 | 22.280 | 15.303 | 0.430 | 21.990 | 17.100 | 0.486 | 22.300 | 20.901 | 0.697 | 22.363 | |
| TiO | 12.914 | 0.414 | 18.230 | 14.581 | 0.468 | 17.880 | 16.184 | 0.637 | 17.720 | 19.852 | 0.827 | 17.903 | |
| E | 24 MeV | 26 MeV | 31 MeV | ||||||||||
| seg | tgt. | flux () | flux () | flux () | |||||||||
| TiO | 24.111 | 0.596 | 23.525 | 26.227 | 0.688 | 24.089 | 31.040 | 0.933 | 24.850 | ||||
| TiO | 22.982 | 0.844 | 21.974 | 24.906 | 0.969 | 22.019 | 29.237 | 1.295 | 22.083 | ||||
| TiO | 21.714 | 0.993 | 17.694 | 23.425 | 1.144 | 17.238 | 27.230 | 1.532 | 16.818 | ||||

### v2

# Photo-nuclear cross sections on 197Au, an update on the gold standard

@@ Abstract
Cross sections for the 197Au reaction are broadly used in nuclear physics as a standard for normalizing photonuclear reaction cross-section data at photon beam energies above approximately 8 MeV. In this paper, we report cross-section measurements for the 197AuAug+m1 reaction at beam energies from 13 to 31 MeV. Our measurements provide the first cross-section data for this reaction at beam energies above 20 MeV, enabling the use of this reaction as a cross-section standard up to 30 MeV. Also, this work provides first cross-section measurements for the 197AuAum2 reaction. In addition, we measured cross-section data for the 197AuAu reaction, which can be used as a cross-section standard above about 25 MeV. These measurements were performed using a new target activation method that is based on the angle-energy correlation of the laser Compton-scattered photon beams at the High Intensity Gamma-ray Source (HIS). The technique enables measuring photonuclear reaction cross-sections at several discrete beam energies concurrently via a single irradiation on a stack of different targets. Measurements were carried out by irradiating a stack of concentric-ring targets consisting of Au, TiO2, Zn, Os, and Au (in order of the -ray beam direction). Our data for the 197Au(,n)196Aug+m1 reaction in the energy range of 13 to 20 MeV are in good agreement with existing ones measured using monoenergetic -ray beams, but differ from data acquired using a bremsstrahlung -ray beam. Also, above 18 MeV, our data for the 197Au(,n)196Aug+m1 and 197Au(,n)196Aum2 reactions differ significantly from the most recent TENDL and JENDL evaluations, suggesting a need to update these data libraries. The TENDL evaluation and existing data are consistent with our data for the 197Au(,3n) reaction, but differ significantly from the JENDL evaluation above 26 MeV.
@@ I Introduction
Radioisotopes play important roles in numerous fields ranging from medical treatments to national security and basic research. Some examples include investigations of structures and reactions involving atomic nuclei, Mossbauer spectroscopy, radio-thermoelectric generation and other nuclear batteries, nuclear device detection, and the mitigation of nuclear proliferation [1]. The reports from “Workshop on the Nation’s Needs for Isotopes: Present and Future”, “Isotopes for the Nation’s Future, A long range plan”, and “Meeting Isotope Needs and Capturing Opportunities for the Future” identify multiple isotopes currently in short supply [1, 2, 3]. Among these are high specific activity beta emitters such as 47Sc, 67Cu, 77As, and 186Re. These radioisotopes are of interest to the community as they have ideal nuclear properties for medical applications. Their availability is limited as they are difficult to produce with high specific activity using common production methods.
@@
Most radioisotopes are produced using either nuclear reactors or light-ion accelerators. High specific activity radioisotopes are very important in nearly all aspects of radiochemical work. Specific activity refers to the radioactivity of the produced radionuclide relative to the total mass of the element of interest present in the sample. Most radionuclides produced in nuclear reactors have low specific activity due to the inherently limited production pathways, e.g., dominated by (n,) reactions. In contrast, radionuclides produced at particle accelerator facilities via charged-particle induced reactions are generally characterized by high specific activity, e.g., (p,xn) reactions are typically used because the produced isotope is a different element from the target material, allowing chemistry techniques for separating materials. Also, photonuclear reactions, e.g., (,x), provide a promising process for producing high specific activity radioisotopes. The Nuclear Science Advisory Committee (NSAC) recently identified the use of photonuclear reactions as one of the most compelling opportunities for large-scale production of high specific activity radioisotopes using high-flux bremsstrahlung -ray beams generated with high-current linear accelerators (LINACs) [1]. The isotope production rate is directly proportional to the integral of the product of the -ray beam flux on the target and the photonuclear cross section. The end-point energy of the bremsstrahlung -ray beam spectrum is usually adjusted to optimize isotope production over the Giant Dipole Resonance (GDR) energy region. Accurate cross-section data for photonuclear reactions over this GDR energy region are needed for developing strategies for radioisotope production with high specific activity. A great deal of work has been done on measuring experimental cross-section databases for some reactions, e.g., (,n) reactions [4, 5, 6]. However, more measurements are needed. For example, the cross-section data for (,p) reactions (which lead to high specific activity radioisotopes) still require considerable work. There are many theoretical databases available which provide guidance on photo-production of radionuclides. Experimental data are needed to evaluate the accuracy of predictions of the leading reaction models used in nuclear-data libraries (TENDL, JENDL, ENDF, JEFF, CENDL, BROND, etc.) [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Examples of theoretically predicted cross sections which incorrectly estimate the production rate of radioisotopes are reported in the literature [17, 18, 19, 20]. Photonuclear cross sections are typically measured using broad energy spectrum bremsstrahlung -ray beams (either untagged or tagged) or quasi-monoenergetic -ray beams produced via laser Compton backscattering (LCB) or positron in-flight annihilation. For quasi-monoenergetic beams, LCB is the modern beam production method, see Table 1. Compared with a bremsstrahlung beam, which has a broadband energy intensity distribution, a LCB -ray beam is narrowly peaked around a specific energy. Precise measurement of cross sections can be performed using activation techniques with target irradiation by such LCB -ray beams. Independent of the type of -ray beam used, cross-section measurements require accurate determination of the beam flux. One approach is to establish a cross-section standard that can be used to determine the beam flux over a specified beam energy range. The cross sections for the 197Au(,xn) reactions are widely used as a standard for determining the photon beam flux in photonuclear reaction measurements. As a standard for cross-section normalization, errors in 197Au(,xn) cross-section data will impact a broad range of nuclear physics research and applications. For example, gold monitor foils are routinely used to normalize bremsstrahlung photon beams in cross-section measurements [21, 22, 23]. The current status of 197Au(,x) cross-section data is as follows. The 197Au(,n)196Au cross section data from different experiments are in good agreement in the energy range of 8-14 MeV, which is below the threshold for multiple neutron emission (see Table 1). However, there are significant discrepancies between data [24, 25] at photon energies between 14-18 MeV. Also, there are no cross-section data for the 197Au(,n) reaction above 20 MeV and extrapolation of data evaluations by TENDL and JENDL differ substantially.
@@
In this paper, we describe our newly developed experimental techniques for measuring photonuclear cross sections at discrete energies using a LCB -ray beam and its application to measuring cross sections for photon-induced reactions on 197Au. The photon energies in this work were chosen to be above the threshold for reactions leading to high specific activity radionuclides via the (,p) reaction for use in medical applications such as 48Ti(,p)47Sc and 68Zn(,p)67Cu at around 20 MeV and higher. The technique is based on the well-defined radial energy dependence of a LCB -ray beam which enables measurement of the corresponding spatially dependent cross section. The central portion of the -ray beam has the highest energy and flux density, with the beam energy and flux density decreasing radially outward from the center in a predictable manner [26]. Activation of concentric ring targets by an uncollimated beam will provide experimental activation data to be used to determine cross sections at the average energy of the beam incident on each ring. This method was developed and used for the first time at the Triangle Universities Nuclear Laboratory’s (TUNL) High-Intensity Gamma-Ray Source (HIS). The HIS facility is described by Weller et al. in Ref. [27]. A goal of the present work is to measure the exclusive cross-section excitation functions of the 197Au(,n) and (,3n) reactions for use as cross-section standards. Table 1 provides a summary of the world data for the 197Au(,xn) reactions by beam type. An interesting feature of the existing data is that there is apparently a strong correlation between the beam type (quasi-monoenergetic or bremsstrahlung) and the experimental technique. All experiments in Table 1 using a quasi-monoenergetic -ray beam measured the cross section with direct neutron counting. The measurements using a bremsstrahlung -ray beam determined the cross section via activation techniques. Our measurements are the first to be done with a monoenergetic -ray beam while using activation techniques. In Section V, our data are compared to the results of the previous measurements listed in Table 1. The results of this work for the Ti, Zn, and Os targets measured concurrently with Au will be presented in a subsequent paper. Here, our multiple target multi-energy technique is described in section II, the HIS LCB beam is presented in section III, the experiment is discussed in section IV, section V describes the cross-section analysis, and the results are presented and discussed in section VI.
@@
| (,n) | O. Itoh et al. [31] | 2011 | 8.08-13.14 | 30 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,n) | F. Kitatani et al. [32] | 2011 | 9.3-12.4 | 12 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,n) | F. Kitatani et al. [33] | 2010 | 10.0-14.6 | 8 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,n) | K.Y. Hara et al. [34] | 2007 | 8.23-12.3 | 24 | AIST, Tsukuba, Japan | LCB source | neutron counting |
@@ II Method
This section provides details about the method applied in this work to measure the exclusive (,n) and (,3n) cross sections via the activation method. The excitation functions reported here cover the -energy range from 13 to 31 MeV. The photonuclear cross sections over the range of the Giant Dipole Resonance (GDR) vary by orders of magnitude. The goal of this work is to demonstrate our activation techniques by applying it in measuring the excitation functions of the 197Au(,xn) reaction cross sections. with an energy resolution of 3-10 at each energy. At HIS, the beam energy spread depends on the diameter of the beam defining collimation. In a typical setup with an energy spread (FWHM) of 5 (FWHM), the flux transmitted through the collimator is about 7.5 of the total flux from the Compton-scattered laser beam. Our method of using three radial target segments enables concurrent measurement of the cross section at three central beam energies. The radial energy and flux distribution of the HIS beam was verified experimentally by blocking portions of the beam with tungsten plugs to mimic the concentric-rings target design. Simulations were performed and compared to the experimental results of these calibrations to determine the accuracy of the beam simulations.
@@ II.1 Target stack
In the activation method, the photon beam induces radioactivity via photon-induced reactions, such as (,n) and (,p), in targets. Reaction yields are determined by gamma counting the targets after activation. The target foils used in these experiments are made of separable concentric rings that exploit the energy spread of the HIS beam and enable the activation of multiple targets across multiple energies in a single beam exposure. A concept of a target stack is illustrated in Figure 1. The relatively low and well-characterized attenuation of high-energy photons by the target materials provides the ability to irradiate multiple targets in a single irradiation. The total thickness of the stack of targets used in these experiments was 2.8-3.3 g/cm2, which resulted in a 13 total attenuation of the photon beam at 14 MeV. The entire target may then be exposed to the beam and each individual concentric ring will interact with a narrow range of beam energies and a calculable fraction of the overall photon flux. The individual rings are analyzed with high-purity germanium detectors (HPGe) to determine activation yields. Thus, using this technique, the cross sections at three energies can be measured in a single irradiation. With the stack of five targets and three segments per target, yields for 15 target segments are obtained for each 12-hour run.
@@ II.2 Activation
Using the activation method (see details in Appendix A), the cross sections on the 197Au targets were obtained. The combined cross section for ground and isomeric state is given by,
@@
The cross section for producing 196Au in the 2nd isomer state is,
The cross section for producing 194Au is determined in the same way as Equation 2,
@@
The photon beams are produced by Compton scattering of photons from relativistic electrons circulating in a storage ring at HIS. In the case of relativistic electrons ( 1) scattering at a small angle (1), the scattered photon energy after collision is given by,
where Eg is the scattered photon energy, and Ep and Ee are the energies of the incident photon and electron. The Lorentz factor of the electron is denoted as . The angle is expressed in terms of the radius of the target (r) and the distance from the collision point to the target position, r/L.
As stated above, for each beam irradiation, reaction cross sections were measured for each target at three average beam energies, one for each target segment. The cross sections were determined from the measured activation of each target segment using Equations 14 – 16 (see Appendix A). The parameter in the equations is the -ray beam flux incident on each target segment, i.e., , where is the total beam flux in the direction of the target and is the fraction of incident on ring i. The uncertainty in , , is the dominant systematic uncertainty in the cross-section measurements. There are two systematic uncertainties contributing to : (1) the uncertainty in the measurement of , and (2) the uncertainty in the calculated value of . The measurements are described in section IV.2. The determination of for each segment and its uncertainty is described in this section. We used the model of Sun and Wu [26] to simulate the Compton-scattered photons passing through the free-electron laser (FEL) aperture systems to obtain the beam intensity profile (i.e., , the beam intensity as a function of the radial distance from the beam axis) and the photon energy spectrum (i.e., , the beam intensity as a function of the photon energy . The calculated ) and from the simulations have been benchmarked with other codes such as CAIN by Yokoya [35] and using experimental measurements at HIS [26]. An example of the Monte-Carlo simulations for our experiment is given in Figure 2. For this simulation, the -ray beam at HIS is produced with 660 MeV and a free electron laser photon wave length nm. The radial extent of each ring is given in Figure 2. The energies of peak intensity of the scattered -ray beam at the target location (L 53 m) incident on the three segments are 18, 17 and 16 MeV, respectively. The beam intensity distributions incident on each ring, and , are plotted in Figure 2 a) and b), respectively. For the beam energy range of 13 to 31 MeV covered in our measurements, the typical percentile energy spreads (1) of the photon beam incident on the , and target rings were 1.3-2.7, 2.2-4.3 and 3.2-5.7 of the central beam energy, respectively. The and distributions were computed from the simulations and used to calculate the values of and , the photon energy of the peak intensity of the distribution for each ring. We performed benchmark measurements to assess the systematic uncertainties in the calculations of and using the simulated data. These measurements were performed using rays produced under conditions identical to those used in the simulations, MeV and nm. Special beam collimator arrangements were used to select sections of the beam cross section like the rings in the targets. Each collimator configuration consisted of a lead collimator with an open circular aperture, or the same lead collimator with a circular tungsten rod insert. The lead collimator and tungsten rods were 6 inches (15.2 cm) in length. The axes of the collimator aperture and the cylindrical rod were centered on the beam axis. At the central beam energy of about 25 MeV, the -ray attenuation through the body of the lead collimators and tungsten rods were about and , respectively. The energy spectrum of the beam was measured using a large cylindrical NaI detector (10” dia. 14” thick.), referred to as “Molly”. The efficiency of Molly is greater than 98 at the energies of these measurements. For the beam energy profile measurements, Molly was positioned with its axis centered on the -ray beam axis, and copper attenuators were inserted in the beam path to reduce the beam flux incident on Molly to less than about , where the pulse-pileup rate is negligible. Measurements of were performed for the first four collimator configurations itemized in Table 2 a): N1 = 0.0 to 6.35 mm; N2 = 0.0 to 15.9 mm; N3 = 6.35 to 15.9 mm; and N4 = 11.1 to 15.9 mm. These measurements were carried out using two collimators with aperture diameters of 0.5” (12.7 mm) and 1.25” (31.8 mm) and two tungsten rods with diameters 0.5” (12.7 mm) and 0.875” (22.2 mm). In addition, the difference in the measured for two collimator configurations was determined by comparing measured data to simulations for radial ranges similar to the target rings, e.g., see configuration 5 in Table 2. The values for obtained from the simulated distributions differ from those determined using the measured distributions by about 2, see the results for collimator configurations 1–4 in Table 2. This uncertainty is significantly smaller than the error in our determination of the absolute beam flux, see section IV.2. As shown in Table 2, the mean beam energy computed using the simulated data is in good agreement with the measured mean beam energy in each segment. Examples of measured distributions for several collimator configurations can be found in Figure 3.
| Coll. | Coll. Dia. | Rod Dia. | Meas. | Sim. | Meas. | Sim. | Note | ||
| Config. | mm | mm | mm | mm | MeV | MeV | |||
| 1 | 12.7 | none | 0.0 | 6.35 | 18 | 20 | 27.1 | 27.1 | |
| 2 | 31.8 | none | 0.0 | 15.9 | 100 | 100 | 25.2 | 25.6 | (a) |
| 3 | 31.8 | 12.7 | 6.35 | 15.9 | 81 | 80 | 24.9 | 25.6 | |
| 4 | 31.8 | 22.2 | 11.1 | 15.9 | 45 | 45 | 24.1 | 24.5 | |
| 5 | 31.8 | 12.7/22.2 | 6.35 | 11.1 | 37 | 35 | 25.8 | 26.1 | (b) |